OpenStudy (mgriffin16):

Find the inverse function of g(x)=x+8/3 ? Show all steps. Please help?

OpenStudy (er.mohd.amir):

let y=x+8/3 solve for x gt inverse.

OpenStudy (lolacole12):

Let's solve for g. gx=x+8/3 Step 1: Divide both sides by x. gx ---=x+8/3 ----- x x g=3x+8 _____ 3x Answer: g=3x+8 ----- 3x